OpenStudy (anonymous):

Solve.x+2 over 15 + 3 minus x over 5 = 2 over 3

OpenStudy (pfenn1):

Is this right?\[\frac{x+2}{15} + \frac{3-x}{5} = \frac 23\]

OpenStudy (anonymous):

A. x = 1 B. x = 0.5 C. x = 2 those are my choices

OpenStudy (pfenn1):

What is the LCD?

OpenStudy (anonymous):

3 or 5

OpenStudy (pfenn1):

Actually, it is both 3 and 5, specifically it is 15. Can you rewrite the expression with the denominator=15?

OpenStudy (anonymous):

No i'm confused.

OpenStudy (pfenn1):

I want to find the LCD (lowest common denominator) or lowest common factor of 3, 5, and 15......15 is the lowest number that can be dived by 3, 5, and 15. right?

OpenStudy (pfenn1):

So we need to rewrite the equation so that all the factors have 15 as their denominator. We can't add fractions with all different denominators. We have to come up with a common denominator. So the new equation will be\[\frac{x+2}{15} + \frac{3(3-x)}{15} = \frac{(5)(2)}{15}\]Of course now we can multiply both sides by 15 to get\[x+2+9-3x=10\]Now combine like terms and solve for x.

OpenStudy (anonymous):

lol